# High Redshift Merger Model for Low Frequency Gravitational Wave Background

### v1

@@ 5.1 Motivation of Early Merger Model
Moreover, JWST is expanding our understanding of the high-redshift scaling relation by probing further and into smaller masses. Recent observations suggest that even less massive SMBHs at high redshift are overmassive compared to local galaxies (Pacucci et al., 2023; Mezcua et al., 2024; Juodžbalis et al., 2024). Based on these observations, we propose a strong redshift dependence in the scaling relation for the additional component of the early merger model. Our parameter choices align with results extrapolated from correlation observed in less massive SMBHs (Pacucci et al., 2023; Mezcua et al., 2024), though it is currently not tightly constrained.
This modification could, in principle, apply to the normal component as well, but we adopt a conservative approach in light of studies suggesting no significant evolution in the scaling relation (Ding et al., 2023). However, the stellar masses of high-redshift quasar host galaxies remain highly uncertain (Yue et al., 2024; Mezcua et al., 2024). Given the rapidly evolving research field, we qualitatively reproduce the observed trend and defer detailed modeling to future studies.
@@ 5.2 Validity of Assumptions
This reasoning can also be reflected in Agazie et al. (2023b), where the best fitting results from GW-only and phenomenological evolution (contains environmental interactions) generate almost identical spectra at higher frequencies. Therefore, the late merger model used in Agazie et al. (2023b) will prefer the same population for galaxy mergers and high-frequency binaries based on the current detected GWB, regardless of the choice of environmental interactions.
Environmental interactions, in this context, serve to bridge binaries to the threshold of GW-driven inspiral, introducing a temporal delay between the mergers of SMBHs and their host galaxies. Since the lifetime of binaries is still poorly understood and not well-constrained (Agazie et al., 2023b), we just assume a constant binary total lifetime of 0.1 Gyr.
Additionally, if a binary has eccentricity, it distributes GW energy across multiple integer harmonics. This redistribution results in a shift of energy towards higher frequencies, manifesting as a low-frequency turnover, a high-frequency flattening of the spectrum, and an intermediate “bump” (Enoki & Nagashima, 2007; Chen et al., 2017). However, the pronounced manifestation of these spectral features requires binaries to have very high eccentricities (e.g., $e\gtrsim 0.9$) while having very small separations, which is unlikely to happen. Therefore, we can safely restrict our analysis to circular binaries only.
@@ 6 Conclusions and Outlook
In light of recent observations, we propose the early merger model, where an additional component of the SMBHs merged at greater redshifts, compared to the standard scenario where almost all the mergers take palce at low redshift. The two models produce similar GWB but exhibit several distinct observational outcomes that can be distinguished in future observations:
@@
Moreover, JWST is detecting more high-redshift AGN than previously expected (Matthee et al., 2024; Greene et al., 2024). Given that galaxy mergers are commonly associated with AGN activity, this provides further evidence in favor of the early merger model, which future observations will need to confirm.
If the early merger model is favored over the standard late merger model by future observations, it will greatly enhance our understanding of galaxy formation and evolution in the early universe. It implies that the evolution pathways of galaxies at the high-mass end of the GSMF might differ significantly from those of less massive galaxies.
@@ acknowledgments
Appendix
Table 1: Summary of model components with their symbols and values.

### v2

@@ 5.1 Motivation of Early Merger Model
[…]
This modification could, in principle, apply to the normal component as well, but we adopt a conservative approach in light of studies suggesting no significant evolution in the scaling relation (Ding et al., 2023). After all, the stellar masses of high-redshift quasar host galaxies remain highly uncertain (Yue et al., 2024; Mezcua et al., 2024). Given the rapidly evolving research field, we qualitatively reproduce the observed trend and defer detailed modeling to future studies.
@@ 5.2 Validity of Assumptions
This reasoning can also be reflected in Agazie et al. (2023b), where the best fitting results from GW-only and phenomenological evolution (contains environmental interactions) generate almost identical spectra at higher frequencies. Therefore, the late merger model used in Agazie et al. (2023b) will prefer the same population for galaxy mergers and high-frequency binaries based on the currently detected GWB, regardless of the choice of environmental interactions.
Environmental interactions, in this context, serve to bridge binaries to the threshold of GW-driven inspiral, introducing a temporal delay between the mergers of SMBHs and their host galaxies, in addition to the dynamical friction timescale. We assume efficient environmental coupling and approximate the total lifetime of binaries using the dynamical friction timescale from Binney & Tremaine (2008). This timescale characterizes the time taken for SMBHs to reach the center of their already merged host galaxies. For a typical velocity dispersion of $\sigma=200\mathrm{~{}km}/\mathrm{s}$ and an initial separation of $r_{i}=5\mathrm{kpc}$ between the SMBHs and the center of the merged galaxy, the timescale is approximately $t_{\mathrm{dy}}\sim 0.1\mathrm{Gyr}$ for SMBHs with a mass of $M\sim 10^{10}M_{\odot}$. Consequently, we approximate the lifetime of binaries to be $0.1\mathrm{Gyr}$ for all mergers.
Additionally, if a binary has eccentricity, it distributes GW energy across multiple integer harmonics. This redistribution results in a shift of energy toward higher frequencies, manifesting as a low-frequency turnover, a high-frequency flattening of the spectrum, and an intermediate “bump” (Enoki & Nagashima, 2007; Chen et al., 2017). Furthermore, binary eccentricity reduces the hardening timescale, especially for low-frequency binaries (Sesana, 2013b; EPTA Collaboration et al., 2024). However, the pronounced manifestation of these spectral features requires binaries to have very high eccentricities (e.g., $e\gtrsim 0.9$) while having very small separations, which is unlikely to happen (Agazie et al., 2023b). Therefore, we can safely restrict our analysis to circular binaries only.
@@ 6 Conclusions and Outlook
In light of recent observations, we propose the early merger model, where an additional component of the SMBHs merged at greater redshifts, compared to the standard scenario where almost all the mergers take place at low redshift. The two models produce similar GWBs but exhibit several distinct observational outcomes that can be distinguished in future observations.
@@
Moreover, JWST is detecting more high-redshift AGNs than previously expected (Matthee et al., 2024; Greene et al., 2024). Given that galaxy mergers are commonly associated with AGN activity, this provides further evidence in favor of the early merger model, which future observations will need to confirm.
If the early merger model is favored over the standard late merger model by future observations, it will greatly enhance our understanding of galaxy formation and evolution in the early Universe. This implies that the evolution pathways of galaxies at the high-mass end of the GSMF might differ significantly from those of less massive galaxies.
@@ acknowledgments
Table 1: Summary of model components with their symbols and values. All parameter values of the GOMV and Late Merger models align with those in Agazie et al. (2023b).